# Braneworld Black Bounce to Transversable Wormhole

### v1

@@ IV.1 Energy conditions
$\rho_{E}+p_{1}$ $=$ $-\frac{2a^{2}z^{2}}{l^{2}\kappa_{5}^{2}}\left(\frac{(r^{2}+a^{2})^{1/2}-2M}{(r^{2}+a^{2})^{5/2}}\right),$
$\rho_{E}+p_{2}$ $=$ $\frac{3a^{2}z^{2}}{l^{2}\kappa_{5}^{2}}\cdot\frac{M}{(r^{2}+a^{2})^{5/2}},$
From the above expressions, it is easy to see that NEC is satisfied for all values of $r$ and $z$ for $\rho_{E}+p_{2}$ and $\rho_{E}+p_{3}$ In Fig. (2) we have the plot of $\rho_{E}+p_{1}$ as a function of $r$ for a fixed value of $z$, where it is possible to observe that this condition is satisfied only in a certain region only for the RBH case, and also the asymptotic radial limit of the expression $\rho_{E}+p_{1}$. Thus, from Equation (37), we can see that NEC is only satisfied in this case for $r\leq\sqrt{(2M)^{2}-a^{2}}$, i.e, $r\leq r_{h}$ for the RBH case and at the limit $r\to\infty$.
For the Strong Energy Conditions (SEC), characterized by $\rho_{E}+\sum_{i}p_{i}\geq 0$, we have that
$\rho_{E}+\sum_{i}p_{i}=\frac{1}{l^{2}\kappa_{5}^{2}}\left(18+\frac{a^{2}z^{2}}{\tilde{r}^{5}}(\tilde{r}-M)\right)$ (38)
In Fig. 3, we have a plot of the above expression, when it’s possible to see that equation (38) is positive when we consider a fixed point in $z$ coordinate. In the right panel of the figure, we can see that, for adequate parameters, the expression can assume negative values.
@@
Thus, the energy conditions SEC are always satisfied in the bulk. While NEC and WEC may or may not be satisfied depending on the value parameters. This result differs from other braneworld models where NEC can be violated in the bulk Neves (2021).
@@
For some parameter values, our results are consistent with observations of the shadows of M87* and the diameter of the image of Sgr A*, which have provided indications that the induced geometry on the brane could represent both ultra-compact objects and wormholes. In particular, reference Banerjee et al. (2022) claims that, even though wormholes, in general, require exotic matter to remain stable, in the presence of extra dimensions such is not the case. Therefore, they showed that the wormhole solution considered does not require exotic matter on the brane.
IV.2 Stability of the solution
@@
We will now investigate the stability of the obtained braneworld black bounce solutions by analyzing the sound velocities of the fluid along radial and extra-dimensional directions. From equations (13) and (36) we get
@@
using the prime to denote the derivative with respect to coordinate $r$ (coordinate $z$). We have used the fact that the average pressure $<p>$ is given by
@@
As clarified at the end of section II, to ensure stability in the solution, we must impose $0\leq(v^{i}_{s})^{2}\leq 1$, or
@@
Since we have most of the regions with instability, the conclusion is that solutions persist to be unstable. However, it is worth noting that we have some stable regions, which is a very different result compared to the original Hawking model.
We finalize by commenting on stability under wavelength perturbations. Now we have $r_{e}=\sqrt{(2M_{e})^{2}-a^{2}}$, with $(2M_{e})^{2}>a^{2}$. Assuming that the AdS radius $l$ is on the order of the Planck length, it should hold that $M/z_{0}>>1$ to be consistent with astrophysical data. In contrast to the non-regularized solution of Chamblin et al. (2000), where the spacetime is AdS everywhere, our regularized solution only exhibits AdS behavior at the asymptotic end of the radial coordinate. In our case, the spacetime is no longer flat for finite values of the radial coordinate and large values of $l$. Thus, it is no longer accurate to claim that the black string solution is unstable at the AdS horizon for small wavelengths. On the other hand, since our regularized solution behaves as AdS only at the asymptotic end of the radial coordinate, we can only assert that, for wavelengths much greater than $l$, the instability is prevented in this boundary. Although defining a proper event horizon is impossible, we can infer that this should occur far from the AdS horizon. Also, at the asymptote of the radial coordinate, the space-time behaves almost flat for wavelengths much smaller than $l$ near the AdS horizon. Thus, this latter location could be unstable. It is worth mentioning that it becomes intriguing to test whether the regularized black string solution is unstable for finite values of the radial coordinate and at the AdS horizon, where the space-time is no longer flat, as in the non-regularized case. This aspect could be studied in future work.
V Conclusion and Summarize
@@
Moreover, both the geometrical and matter deformations, represented by equations (18) and (19) respectively, result in the induced geometry on the brane transitioning from a singular Schwarzschild spacetime to a regularized SV spacetime. Given the absence of sources or matter fields on the brane, we can assert that the induced SV geometry on the brane arises from the influence of geometrical and matter deformations in the bulk. In this context, the $5D$ spacetime is regular at the origin $r=0$ (or $\tilde{r}=a$), unlike Chamblin et al. (2000), where singularity occurs at the origin. However, the singularity at the AdS horizon persists. This latter singularity could be suppressed in a scenario where the extra dimension is compactified.
Despite being asymptotically AdS, the above deformation induces a transition from a vacuum distribution to an anisotropic distribution. This behavior differs from the black string of Hawking et.al, which is globally AdSChamblin et al. (2000). In fact, by studying the energy conditions, we found that the NEC and SEC are always satisfied in bulk, while the WEC may or may not be satisfied depending on the values of the parameters. This result differs from other Brane world models, where the NEC can be violated in bulk Neves (2021). An interesting result is that, for some parameter values, our results are consistent with observations of the shadows of M87* and the diameter of the image of Sgr A*, which have provided indications that the induced geometry on the brane could represent both ultra-compact objects and wormholes. In particular, reference Banerjee et al. (2022) claims that even though wormholes generally require exotic matter to remain stable, such is not the case in the presence of extra dimensions. Therefore, they showed that the wormhole solution considered does not require exotic matter on the brane.
@@
We have also scrutinized the stability (and physical validity) of the braneworld black-bounce solutions in bulk by evaluating the sound speed of the source along the $r$ and $z$ directions. Our analysis unveiled specific regions where stability and physicality are guaranteed for both wormhole and regular black hole solutions, contingent upon the bounce parameter $a$, as shown in Fig. 4. It is worth noting that in braneworld models where $a=0$, there are no stable regions. Because the spacetime is only AdS at the asymptote of the radial coordinate, we can assert that only at that location the solution is unstable considering now wavelength perturbations near the AdS horizon but stable far from it. Thus, it becomes intriguing to test whether our solution is unstable for finite values of the radial coordinate, where the spacetime is no longer AdS as the black string of reference Chamblin et al. (2000).
@@
Finally, we mention that in the case of the coordinate $z$, we cannot do a Simpson-Visser-like regularization $z\to\sqrt{z^{2}+b^{2}}$, as we would continue to have a divergence in $z\to\infty$. An alternative would be to consider an exchange of the type $z\to f_{b}(z)$ so that $\lim_{z\to\infty}f_{b}(z)$ is a constant and $\lim_{b\to 0}f_{b}(z)=z$. However, nothing guarantees that such a function exists, and if it does, imposing an analytical form for $f_{b}(z)$ does not seem to be trivial. It is also necessary to test whether the regularized black string solution is unstable for finite values of the radial coordinate and at the AdS horizon, where the space-time is no longer flat, as in the non-regularized case. These aspects will explored in future work.

### v2

@@ III.1.3 SEC
As mentioned in the introduction, observations of the shadows of M87* and the diameter of the image of Sgr A* have provided indications that the induced geometry on the brane could represent both ultra-compact objects and wormholes. Specifically, reference Banerjee et al. (2022) argues that, although wormholes generally require exotic matter to remain stable, this is not necessarily the case in the presence of extra dimensions. They demonstrated that the wormhole solution they considered does not require exotic matter on the brane. Similarly, in our case, no additional exotic matter is needed on the brane beyond its own tension $\sigma$. However, since the energy conditions are not satisfied in several scenarios within the bulk, the inclusion of exotic matter in the bulk is necessary to sustain the induced geometry on the brane.
III.2 Stability of the solution
III.2.1 Discussion about wavelength perturbations and our geometry
It is well known that black string solutions in asymptotically flat space are unstable to wavelength perturbations Gregory and Laflamme (1993). As noted by Chamblin et al. (2000), an AdS spacetime acts as a confining box, preventing the development of fluctuations with wavelengths much larger than $l$, thereby mitigating this type of instability. In this connection, if the curvature scalar $R=-20/l^{2}$ is very small, the AdS curvature becomes negligible, and the spacetime behaves like flat space. This implies that the black string solution will be unstable to perturbations with wavelengths on the order of the horizon radius $r_{e}=\frac{2Ml}{z_{0}}$, such that $\frac{r_{e}}{l}=\frac{2M}{z_{0}}\ll 1$. In other words, for very large values of $z$, such perturbations will fit within the AdS box, and instability could occur near the AdS horizon. However, for small values of $z_{0}$, potential instability would occur at wavelengths much larger than $l$, i.e., $\frac{r_{e}}{l}=\frac{2M}{z_{0}}\gg 1$, and thus the AdS curvature prevents instability. Therefore, the black string solution is stable far from the AdS horizon but potentially unstable near it. This issue was addressed when it was shown that RS black holes are unstable under linear perturbations Gregory (2000).
In contrast to the previous paragraph, our regularized solution only exhibits AdS behavior at the asymptotic end of the radial coordinate. Thus, our spacetime is not nearly flat for large values of $l$ at finite radial coordinate values, as is the case in an AdS black string. Therefore, it is not accurate to claim that our solution is unstable near the AdS horizon and stable far from it. Such assertions would only be valid at the radial coordinate boundary where our geometry behaves like an AdS space. This makes it intriguing to study the stability under wavelength perturbations for finite radial coordinate values. However, this investigation extends beyond the scope of our current research and could be explored in future work.
III.2.2 Speed of sound criteria
The fact that the criterion described in the previous section appears to be incompatible with our geometry prompts us to consider alternative criteria. In this context, it is well known that the stability criterion and causality associated with the speed of sound have been applied in the literature to solutions featuring matter in the energy-momentum tensor Capozziello and Godani (2022); Das et al. (2023). Specifically, the speed of sound in a direction $i$ is defined as:
$(v^{i}_{s})^{2}=\frac{dp_{i}}{d\rho},$ (35)
See, for example, Appendix II-A of Reference Armendariz-Picon and Lim (2005), where the last formula is derived from perturbations of the Lagrangian dependent on the squared gradient of a scalar field. Also, refer to Chapter 1.7 of Reference Vietri (2008), where the previous formula is derived by applying perturbations to the density and pressure in the mass conservation equation.
The stability condition states that $0\leq(v^{i}_{s})^{2}\leq 1$ Capozziello and Godani (2022); Das et al. (2023). Reference Ellis et al. (2007) indicates that the situation where $(v^{i}_{s})^{2}>1$ implies that the speed of sound cones lie outside the light cones in all directions at all events. Consequently, fluid waves can propagate at speeds up to and including this superluminal speed of sound. This reference also notes that such a situation is far from what is typically observed in ordinary matter.
First of all, we can observe that at infinity, the equation of state is recovered as $p_{i}\sim-\Lambda\sim\rho$. This is because, at infinity, the space exhibits an AdS (Anti-de Sitter) geometry. Consequently, in the parameter space, we can deduce that $\frac{d<p>}{d\rho}=-1$. As noted in reference Giambò and Luongo (2024), it is important to emphasize that the sound speed represents the velocity of perturbations in a given fluid. In the case of negative equation of state parameters, it is possible for the squared sound speed to become negative, leading to instabilities.
@@ III.2.2 Speed of sound criteria
Thus, in the context of braneworld models with the presence of matter fields in the bulk, we propose to study stability based on the causality criterion of the speed of sound. To do this, we will use the assumption that all pressures propagate uniformly in all spatial directions of the bulk. In this way, We will now investigate the stability of the obtained braneworld black bounce solutions by analyzing the sound velocities of the fluid along radial and extra-dimensional directions. From equations (35) and (26) we get
@@
in the coordinates space, using the prime to denote the derivative concerning coordinate $r$ (coordinate $z$). We have used the fact that the average pressure $<p>$ is given by
@@
As was mentioned, to ensure stability in the solution, we must impose $0\leq(v^{i}_{s})^{2}\leq 1$, or
@@
Since we have most of the regions with instability, the conclusion is that solutions persist to be unstable. However, it is worth noting that we have some stable regions, which is a very different result compared to the vacuum AdS case.
It is important to note that, as might be expected, instabilities also occur in the coordinate space. This can be easily observed by taking the limits $\lim_{r\to\infty}(v^{r}_{s})^{2}=\lim_{r\to\infty}(v^{z}_{s})^{2}=-\frac{1}{2}$. This is expected because, at the boundary, the geometry behaves like an AdS (Anti-de Sitter) space.
IV Conclusion and Summarize
@@
Moreover, both the geometrical and matter deformations, represented by equations (5) and (6) respectively, result in the induced geometry on the brane transitioning from a singular Schwarzschild spacetime to a regularized SV spacetime. Given the absence of sources or matter fields on the brane, we can assert that the induced SV geometry on the brane arises from the influence of geometrical and matter deformations in the bulk. In this context, the $5D$ spacetime is regular at the origin $r=0$ (or $\tilde{r}=a$), unlike the AdS case, where singularity occurs at the origin. However, the singularity at the AdS horizon persists. This latter singularity could be suppressed in a scenario where the extra dimension is compactified. Despite being asymptotically AdS, the above deformation induces a transition from a vacuum distribution to an anisotropic distribution.
Our study of the energy conditions reveals that the Strong Energy Condition (SEC) may or may not be satisfied depending on the parameter values. From the perspective of the brane itself, our results are consistent with observations of the shadows of M87* and the diameter of the image of Sgr A*. These observations suggest that the induced geometry on the brane could represent both ultra-compact objects and wormholes. Specifically, reference Banerjee et al. (2022) argues that while wormholes generally require exotic matter to remain stable, this is not necessary in the presence of extra dimensions because exotic matter is not required on the brane. Similarly, in our case, no additional exotic matter is needed on the brane beyond its own tension $\sigma$. However, since the energy conditions are not satisfied in several scenarios within the bulk, the inclusion of exotic matter in the bulk is necessary to sustain the induced geometry on the brane.
@@
On the other hand, since the spacetime is only AdS at the asymptote of the radial coordinate, we can conclude that the solution is unstable near the AdS horizon but stable far from it when considering wavelength perturbations criterion. This observation prompts us to investigate whether the solution remains stable at finite values of the radial coordinate, where the spacetime deviates from AdS. Consequently, we need to explore alternative criteria for assessing the stability of the geometry at these finite radial distances. In the context of braneworld models with matter fields in the bulk, we propose evaluating stability based on the causality criterion of the speed of sound. We have also analyzed the stability and physical validity of the braneworld black-bounce solutions by examining the sound speed along the $r$ and $z$ directions. Our analysis identified specific regions where stability and physicality are assured for both wormhole and regular black hole solutions, depending on the bounce parameter $a$, as illustrated in Fig. 4. It is important to note that for braneworld models where $a=0$, no stable regions are present. This aligns with the wavelength perturbation criterion because, in the limit $a\to 0$, the spacetime behaves as AdS. These instabilities are also observed in the limit $r\to\infty$, where the spacetime again behaves as AdS.
@@
Finally, we mention that in the case of the coordinate $z$, we cannot do a Simpson-Visser-like regularization $z\to\sqrt{z^{2}+b^{2}}$, as we would continue to have a divergence in $z\to\infty$. An alternative would be to consider an exchange of the type $z\to f_{b}(z)$ so that $\lim_{z\to\infty}f_{b}(z)$ is a constant and $\lim_{b\to 0}f_{b}(z)=z$. However, nothing guarantees that such a function exists, and if it does, imposing an analytical form for $f_{b}(z)$ does not seem to be trivial. It is also necessary to test whether the regularized black string solution is unstable for finite values of the radial coordinate and at the AdS horizon, where the space-time is no longer flat, as in the non-regularized case. These aspects will be explored in future work.
Acknowledgements
Milko Estrada is funded by ANID , FONDECYT de Iniciación en Investigación 2023, Folio 11230247. Tiago M. Crispim, Geová Alencar, and Celio R. Muniz would like to thank Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação Cearense de Apoio ao Desenvolvimento Científico e Tecnológico (FUNCAP) for the financial support.